OpenStudy (anonymous):

I have the answer can someone please check me..... Find the derivative of h(x)=arctan(1/x)

OpenStudy (goformit100):

h(x)=arctan(1/x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got for the answer\[\frac{ x^2 }{ x^2+1}\] is that correct

OpenStudy (goformit100):

Yes it's correct.

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

keep in mind that the answer is negative though

jimthompson5910 (jim_thompson5910):

since the derivative of the inner function 1/x is -1/(x^2)

OpenStudy (anonymous):

wait I'm confused now then what should the answer look like

jimthompson5910 (jim_thompson5910):

h(x) = arctan(1/x) h ' (x) = d/dx [ arctan(1/x) ] h ' (x) = 1/( (1/x)^2 + 1) * d/dx[ 1/x ] h ' (x) = 1/( 1/x^2 + 1) * (-1/x^2) h ' h(x) = (1/x^2)/( 1/x^2 + 1) h ' (x) = -1/(1 + x^2) ... multiply every term by x^2 to clear out the fractions

jimthompson5910 (jim_thompson5910):

So the answer is h ' (x) = -1/(1 + x^2) or h ' (x) = -1/(x^2 + 1)

OpenStudy (anonymous):

ok got it thanks

jimthompson5910 (jim_thompson5910):

yw